Convert propane into 1-paopanol, 2-propanol and allyl alcohol.

Grade:11

1 Answers

Jessica
6 years ago
CH3-CH2-CH3 + Br2=CH3-CH-Br-CH3CH3-CH-Br-CH3 + Koh= CH3-CH=CH2Treating propane with Br gas and then the resultant 2-bromo propane is treated with alcoholic Koh